### v1

@@ 1 Introduction
The General Data Protection Regulation (GDPR) establishes anonymization as a mechanism through which personal data can be rendered outside the scope of data protection laws. Recital 26 of the GDPR defines anonymization as the process by which data is “rendered anonymous in such a manner that the data subject is not or no longer identifiable.” However, the absence of clear, practical guidance on how to achieve this standard has resulted in inconsistent implementations and legal uncertainties. Many organizations either overestimate the effectiveness of their anonymization processes or struggle to comply due to conflicting regulatory interpretations. Additionally, courts have recognized that anonymization is never absolute — what is considered anonymous today may become identifiable tomorrow as technology advances. Despite the importance of anonymization, the regulatory and educational landscape remains fragmented and inadequate. On one end of the spectrum, legal guidelines provide high-level definitions and compliance requirements but lack technical specificity. On the other end, academic research offers rigorous, mathematically grounded approaches to anonymization that are often inaccessible to practitioners who do not have advanced expertise in statistics or computer science. This disconnect has left engineers, data scientists, and policymakers without the necessary tools to implement anonymization effectively. The result is widespread reliance on outdated or insufficient methods—such as k-anonymity and l-diversity—that have been repeatedly shown to fail against modern re-identification attacks [46].
Furthermore, anonymity-washing is exacerbated by inconsistent regulatory interpretations across jurisdictions. The European Union has exercised significant global influence on data privacy regulation, with many countries modelling their laws after the GDPR. However, even within the EU, national data protection authorities and courts have issued conflicting opinions on what constitutes effective anonymization, leading to uncertainty among organizations attempting to comply. Beyond Europe, frameworks such as the United States’ de-identification standards under the Health Insurance Portability and Accountability Act (HIPAA) and the California Consumer Privacy Act (CCPA), Japan’s Act on the Protection of Personal Information (APPI), and emerging guidelines such as the Brazilian General Data Protection Law (LGPD) further demonstrate that approaches to anonymization lack uniformity at the international level, making cross-border data governance exceedingly complex.
@@
First, in Section 2 we introduce the concept of anonymity-washing and situate it within the broader landscape of privacy discourse. In Section 3, we examine the legal foundations and the regulatory ambiguity surrounding anonymization. Next, Section 4 presents an overview and critique of technical guidelines and educational resources, highlighting the gaps practitioners face. Section 5.1 explores the practical implications of anonymity-washing, including legal cases and implementation failures. Finally, Section 6 offers recommendations and concluding reflections on how to address the risks of anonymity-washing through clearer guidance and improved institutional coordination.
@@ 2.1 Anonymization terminology
The question of intentionality behind deceitful anonymity statements deserves a brief focus, as the term “washing” implies an intentional action. Except that intention in this context can be difficult to prove. Sometimes, anonymity-washing cases are so blatant that the willingness to deceive leaves no doubt. Other times, anonymity-washing is harder to prove and therefore appears incidental, giving the impression that data controllers and/or processors are acting in good faith while deploying weaker solutions. There is, of course, a risk of mischaracterization. Still, it may never be possible to prove with a high degree of confidence that a data controller and/or processor acted in good faith, since defendants are likely to claim to be acting in good faith when notified, and in the course of legal proceedings.
@@ 3 Overview of regulatory guidance on data anonymization
On 25 July 2024, the European Commission published its second report on the implementation of the GDPR [44]. One of the key issues highlighted in the report is the persistence of differing interpretations among national data protection authorities, which undermines the uniform application of the GDPR. This discrepancy gives rise to legal uncertainty, thus businesses are confronted with divergent administrative requirements across different Member States. In this regard, the Commission seeks to reiterate its request, previously made in 2020 [35], to support practitioners by providing clearer guidance and materials to facilitate GDPR compliance. This issue is particularly pertinent in the context of anonymity washing, as the Commission has reported in [36]:
@@ 3.1 EU regulations
More recent guidelines have been adopted by the EDPB on issues related to data anonymization, but they do not contain additional advice. On 17 December 2024, the EDPB published guidelines on the anonymization of AI models [38]. It states that whenever models are trained on personal data, they cannot be considered anonymous. The reason is that many studies on these models have demonstrated their capacity to “regurgitate” part of their training datasets [6]. The EDPB stated that a model is considered anonymous only when, based on appropriate documentation, personal data cannot be inferred either directly (by statistical inference, including the probabilistic functioning of the model) or indirectly (within a user’s prompt). If the risk of “regurgitation” of personal data persists, a deeper analysis is needed [^8] The following aspects are required for verification: the source of the data, their preparation and minimization, the training method, the analysis of the model, the resistance of the model to cyber attacks and the documentation provided..
@@ 3.2 Anonymization regimes beyond the EU
Anonymization guidelines are present in data protection regimes across the globe. There are differences, however, in the approaches and the overall granularity levels exhibited by the relevant frameworks. Notably, some data protection regimes, such as in Japan [58] and the US [93], come with relatively detailed guidance on how to achieve the expected levels of anonymization and how to handle the data [85, 5]. With regard to data transfer between the EU and the USA the previous Privacy Shield, which was adopted on the basis of the European Commission’s decision that the USA’s level of personal data protection was equivalent to that of the EU,  [42], was replaced by a revised Privacy Framework, after the EU Court of Justice overturned it [60] In contrast, other regulatory frameworks, such as in Brazil[14], are not particularly prescriptive and require additional input. At the same time, several jurisdictions have initiated efforts to modernize their approaches to data protection, including anonymization. For example, Brazil’s national data protection authority, the ANDP, is set to clarify what measures could be implemented to ensure anonymity in accordance with its 2018 Lei Geral de Proteçao de Dados (LGPD) in the upcoming years [4]. A call for public participation in that effect has been published in early 2024. In 2023, the Data Security Council of India (DSCI) published a roadmap considering possible orientations for a national data anonymization regime [53]. At the intra-state level, in Québec, the Regulation respecting the anonymization of personal information was published in 2024 [82]. The text is very prescriptive and seeks to clarify the distinction between anonymity and pseudonymity, aligning with the EU’s view. In the EU, the EDPB is expected to publish new guidelines on anonymization later this year. It is expected that the new guidelines will fix the inconsistencies introduced by the WP29 Opinion 5/2014 on anonymization, thereby clarifying the dominant approach at the EU level [40].
@@ 4.1 Contradictory guidelines and uncertain standards within the EU
Some authors consider that the EU Data Protection Law lacks a clear definition of anonymization [16]. Unlike pseudonymization, the GDPR fails to define anonymization in its Article 4 titled“definitions”. However, another Opinion of WP29 on the concept of personal data issued in 2007 [79] clarified the difference between anonymization and pseudonymization. Recalling ISO’s previous definitions, they explained that anonymization protects privacy, while pseudonymization represents a technical, reversible process. Nevertheless, data can still be considered anonymous, even when re-identification remains possible, but complementary measures to prevent re-identification are implemented. This flexible approach was not supported in the Opinion 5/2014 on anonymization [78], which applies together with the previous Opinion on the concept of personal data. In Opinion 5/2014, WP29 required the aggregation of data (into group statistics) and the destruction of raw data (identifiers) to ensure correct anonymization. Nevertheless, the objective still remained to prove that the likelihood of re-identification was negligible. On this point, the Commission’s guidelines [34] on the free flow of non-personal data suggest that it is often difficult to assess the effectiveness of an anonymization procedure. Indeed, besides many academic papers [3, 46, 29, 70, 1, 28], even a study commissioned by the European Parliament’s ITRE Committee has shown that it is possible to re-identify supposedly anonymized data [45].
Moreover, national DPAs disagree on how to implement anonymization. The French DPA, CNIL, adopts the WP29’s approach to anonymization; other national DPAs are more flexible. For example, the UK’s ICO (Information Commissioner’s Office) states that [50]:
@@
“Organisations don’t have to be able to prove that it is impossible for any data subject to be identified in order for an anonymization technique to be considered successful. Rather, if it can be shown that it is unlikely that a data subject will be identified given the circumstances of the individual case and the state of technology, the data can be considered anonymous”.
The CNIL’s guidelines on anonymization are of particular interest, as they establish a strict standard to determine whether data can be considered anonymous. They assert that data is anonymous only when it is impossible to re-identify the data subject. However, they recognise that when the risks of singling out, linkability, and inference are not met, data can be deemed anonymous if a subsequent analysis indicates a negligible risk of re-identification [21]. Unfortunately, the definition of what constitutes a negligible risk remains ambiguous.
For these reasons, many controllers have given up on anonymization and prefer pseudonymization [16] even if pseudonymized data remain subject to the GDPR. The problem in this case is that “pseudonymized data are considered personal data, regardless of whether they are, or ever will be, in the hands of a person who holds the key needed for re-identification” [81]. The application of personal data regulation can be challenging, especially in the context of rapid data access, as evidenced by the case of health data (see also Section 5.1).
Take-away The lack of a clear, harmonized definition of anonymization across EU legal texts and among national authorities creates confusion for practitioners. This leads many to favour pseudonymization, even though it offers weaker privacy protection, with data remaining fully subject to GDPR. This regulatory ambiguity undermines the consistent implementation of and weakens trust in anonymization as a reliable privacy safeguard.
@@ 4.2 Technical Documents
  • DPC (Ireland) stands out with well-structured and clear guidance [31, 32] on legal questions, however, it does not offer practical advice.
@@
Some statistical agencies provide additional resources. The National Institute of Statistics and Economic Studies in France (INSEE) offers slides and working papers [54, 56], but most lack practical detail. A notable exception is Bergeat’s work [57] comparing and explaining experiments done by two anonymization software tools: $\mu$-Argus and SDCMicro. It also gives plenty of citations, however, only to sources on statistics (no computer science references). It is aimed at statisticians, and it could serve as a continuation to other introductory materials, but the reader who already has at least a Bachelor’s degree in mathematics or statistics. Other guides, such as [55], focus on confidentiality rules within the French statistical service rather than techniques. Statistical documents only mention statistical tools and use a different language from that of computer scientists. This could, unfortunately, result in ignoring some state-of-the-art methods, such as differential privacy. A good example is the working paper [30] that details the anonymization process applied to a large French administrative database where the authors experimented with different methods including k-anonymity, all-m anonymity, and l-diversity. They mention that they have tried to apply DP, however, they abandoned the experiment due to a lack of expertise.
@@ 5 Inadequate Practices
5.1 Confusions arising from the definition of personal data
@@
EU case law lacks clarity with regard to anonymization practices. The most relevant cases focus on clarifying the concept of personal data. However, the interpretations provided help to assess what an anonymized dataset is not. Moreover, the Court’s application of Recital 26 to real cases provides valuable insight into the question of whether data remains anonymous despite a residual risk of re-identification. In this regard, the General Court’s SRB vs EDPS decision is a good example [43]. SRB (Single Resolution Board) carried out an insolvency procedure against Banco Popular. Within this procedure, some data were processed to assess the eligibility of the participants for compensation. Each participant was identifiable by means of an alphanumeric code generated randomly. The staff processing this data only had access to these codes and not to the key identifying them. The EDPS (European Data Protection Supervisor) considered these data pseudonymized [77], but its decision was challenged before the General Court. Using a risk-based approach, the General Court decided that the data were anonymous. Indeed, according to the Breyer Court of Justice case law[61], the additional information (the key) needed to re-identify the data subjects remained inaccessible to the processing staff. The fact that the staff could not legally access the complementary data that would allow re-identification proved enough to consider that no reasonable means existed to re-identify the data, which thus remained non-personal [43].
5.2 EUCJ’s case-law on personal data
@@
This example shows that the EUCJ’s case law on personal data builds upon its precedents rather than undergoing a radical evolution. On this point, Breyer’s decision [61] was about the dynamic nature of IP addresses. These IP addresses are subject to change with each connection. The plaintiff initiated legal proceedings against the Republic of Germany for its practices concerning the storage and registration of this data. The Court of Justice had to determine whether dynamic IP addresses should be considered personal data for the service provider. The Court decided that the retention of all information by a single individual was not a prerequisite for data being considered as personal. This meant that a third party could retain such re-identifying information, and that this circumstance did not affect the qualification of the data. However, the Court acknowledged that an assessment was necessary to determine the reasonableness of combining this information, taking into account the effort, time, and cost associated with the operation, as well as the accessibility of this additional information (enabling user identification) to the service provider. Given the legal restrictions on such access in Germany, the Court determined that in the absence of legal means to obtain this information, the data in question were not deemed personal. The doctrine posits that two fundamental elements have been applied since the Breyer decision to ascertain the personal nature of data. These are: (1) the distinguishability of the data, defined as the capacity of the data points to identify an individual, and (2) the availability of additional data to “situationally relevant entities” who are capable of associating these data with a physical person[89].
@@
The Scania decision perfectly [64] illustrates this methodology. In this case, the Court was asked to determine the legal status of a vehicle identification number (VIN), a unique alphanumeric code assigned by manufacturers to identify the proprietor of a vehicle. In its decision, the Court stated that the VIN can be personal data for independent operators and vehicle manufacturers if the former have the additional data that enable re-identification, and for the latter if they make the VIN available. The availability of data is considered in conjunction with the capacity of isolating proprietors of the vehicles or all other people who have a title on them. Some authors have suggested an evolution in the interpretation of personal data, attributing this change to the Court’s categorization of independent operators and manufacturers as “situationally relevant entities”, capable of associating VIN with additional identifying information.
@@
Take-away The consistency of the Court’s jurisprudence on the concept of personal data is paramount to contrast the phenomenon of anonymity-washing. This is particularly crucial given the occurrence of poor anonymization practices, which the EUCJ is entitled to sanction.
5.3 Confusion at the institutional level

### v2

@@ 1 Introduction
The General Data Protection Regulation (GDPR) establishes anonymization as a mechanism through which personal data can be rendered outside the scope of data protection laws. Recital 26 of the GDPR defines anonymization as the process by which data is “rendered anonymous in such a manner that the data subject is not or no longer identifiable.” However, the absence of clear, practical guidance on how to achieve this standard has resulted in inconsistent implementations and legal uncertainties. Many organizations either overestimate the effectiveness of their anonymization processes or struggle to comply due to conflicting regulatory interpretations. Additionally, courts have recognized that anonymization is never absolute — what is considered anonymous [^2] In this paper we use the word anonymous whenever we cite or reference a text that has used the same expression. today may become identifiable tomorrow as technology advances. Despite the importance of anonymization, the regulatory and educational landscape remains fragmented and inadequate. On one end of the spectrum, legal guidelines provide high-level definitions and compliance requirements but lack technical specificity. On the other end, academic research offers rigorous, mathematically grounded approaches to anonymization that are often inaccessible to practitioners who do not have advanced expertise in statistics or computer science. This disconnect has left engineers, data scientists, and policymakers without the necessary tools to implement anonymization effectively. The result is widespread reliance on outdated or insufficient methods—such as k-anonymity and l-diversity—that have been repeatedly shown to fail against modern re-identification attacks [56].
Furthermore, anonymity-washing is exacerbated by inconsistent regulatory interpretations across jurisdictions. The European Union has exercised significant global influence on data privacy regulation, with many countries modelling their laws after the GDPR. However, even within the EU, national data protection authorities and courts have issued conflicting opinions on what constitutes effective anonymization, leading to uncertainty among organizations attempting to comply. Beyond Europe, frameworks such as the United States’ de-identification standards under the Health Insurance Portability and Accountability Act (HIPAA) and the California Consumer Privacy Act (CCPA), Japan’s Act on the Protection of Personal Information (APPI), and emerging guidelines such as the Brazilian General Data Protection Law, the Lei Geral de Proteçao de Dados (LGPD), further demonstrate that approaches to anonymization lack uniformity at the international level, making cross-border data governance exceedingly complex.
@@
First, in Section 2 we introduce the concept of anonymity-washing and situate it within the broader landscape of privacy discourse. In Section 3, we examine the legal foundations and the regulatory ambiguity surrounding anonymization. Next, Section 4 presents an overview and critique of technical guidelines and educational resources, highlighting the gaps practitioners face. Section 5 explores the practical implications of anonymity-washing, including legal cases and implementation failures. Finally, Section 6 offers recommendations and concluding reflections on how to address the risks of anonymity-washing through clearer guidance and improved institutional coordination.
@@ 2.1 Anonymization terminology
The question of intentionality behind deceitful anonymity statements deserves a brief focus, as the term “washing” implies an intentional action, motivated by a paucity of resources, time constraints, or uncertain goals regarding projected uses of data records. Except that intention in this context can be difficult to prove. Sometimes, anonymity-washing cases are so blatant that the willingness to deceive leaves no doubt. Other times, anonymity-washing is harder to prove and therefore appears incidental, giving the impression that data controllers and/or processors are acting in good faith while deploying weaker solutions. There is, of course, a risk of mischaracterization. Still, it may never be possible to prove with a high degree of confidence that a data controller and/or processor acted in good faith, since defendants are likely to claim to be acting in good faith when notified, and in the course of legal proceedings.
@@ 3 Overview of regulatory guidance on data anonymization
On 25 July 2024, the European Commission published its second report on the implementation of the GDPR [45]. One of the key issues highlighted in the report is the persistence of differing interpretations among national data protection authorities, which undermines the uniform application of the GDPR. This discrepancy gives rise to legal uncertainty; thus, businesses are confronted with divergent administrative requirements across different Member States. In this regard, the Commission seeks to reiterate its request, previously made in 2020 [41], to support practitioners by providing clearer guidance and materials to facilitate GDPR compliance. This issue is particularly pertinent in the context of anonymity washing, as the Commission has reported in [42]:
@@ 3.1 EU regulations
More recent guidelines have been adopted by the EDPB on issues related to data anonymization, but they do not contain additional advice. On 17 December 2024, the EDPB published guidelines on the anonymization of AI models [43]. It states that whenever models are trained on personal data, they cannot be considered anonymous. The reason is that many studies on these models have demonstrated their capacity to “regurgitate” part of their training datasets [8]. The EDPB stated that a model is considered anonymous only when, based on appropriate documentation, personal data cannot be inferred either directly (by statistical inference, including the probabilistic functioning of the model) or indirectly (within a user’s prompt). If the risk of “regurgitation” of personal data persists, a deeper analysis is needed [^9] The following aspects are required for verification: the source of the data, their preparation and minimization, the training method, the analysis of the model, the resistance of the model to cyber attacks, and the documentation provided..
@@ 3.2 Anonymization regimes beyond the EU
Anonymization guidelines are present in data protection regimes across the globe. There are differences, however, in the approaches and the overall granularity levels exhibited by the relevant frameworks. Notably, some data protection regimes, such as in Japan [66] and the US [93], come with relatively detailed guidance on how to achieve the expected levels of anonymization and how to handle the data [81, 7]. Concerning data transfer between the EU and the USA, the previous Privacy Shield, which was adopted on the basis of the European Commission’s decision that the USA’s level of personal data protection was equivalent to that of the EU,  [38], was replaced by a revised Privacy Framework, after the EU Court of Justice overturned it [49] In contrast, other regulatory frameworks, such as in Brazil[16], are not particularly prescriptive and require additional input. At the same time, several jurisdictions have initiated efforts to modernize their approaches to data protection, including anonymization. For example, Brazil’s national data protection authority, the ANDP, is set to clarify what measures could be implemented to ensure anonymity in accordance with its LGPD in the upcoming years [4]. A call for public participation in that effect has been published in early 2024. In 2023, the Data Security Council of India (DSCI) published a roadmap considering possible orientations for a national data anonymization regime [29]. At the intra-state level, in Québec, the Regulation respecting the anonymization of personal information was published in 2024 [83]. The text is very prescriptive and seeks to clarify the distinction between anonymity and pseudonymity, aligning with the EU’s view. In the EU, the EDPB is expected to publish new guidelines on anonymization later this year. It is expected that the new guidelines will fix the inconsistencies introduced by the WP29 Opinion 5/2014 on anonymization, thereby clarifying the dominant approach at the EU level [36].
@@ 4.1 Contradictory guidelines and uncertain standards within the EU
Some authors consider that the EU Data Protection Law lacks a clear definition of anonymization [18]. Unlike pseudonymization, the GDPR fails to define anonymization in its Article 4 titled“definitions”. The reason is that anonymization and pseudonymization are considered in a binary way [^10] https://techcrunch.com/2017/10/07/how-anonymous-wifi-data-can-still-be-a-privacy-risk/. Indeed, anonymized data is excluded from the scope of the GDPR, while pseudonymized data is entirely subject to it. This is unlike some legislation (such as in Japan), which suggests a lighter regime for pseudonymized data while still imposing obligations on anonymized data. The EU approach fails to acknowledge that there is always a risk of re-identification with anonymisation, and that pseudonymized data may require lighter protection than ’classic’ personal data. This is despite WP 29 acknowledging this risk. On this point, another Opinion of WP29 on the concept of personal data issued in 2007 [6] clarified the difference between anonymization and pseudonymization. Recalling ISO’s previous definitions, they explained that anonymization protects privacy, while pseudonymization represents a technical, reversible process. Nevertheless, data can still be considered anonymous, even when re-identification remains possible, but complementary measures to prevent re-identification are implemented. This flexible approach was not supported in the Opinion 5/2014 on anonymization [5], which applies together with the previous Opinion on the concept of personal data. In Opinion 5/2014, WP29 required the aggregation of data (into group statistics) and the destruction of raw data (identifiers) to ensure correct anonymization. Nevertheless, the objective still remained to prove that the likelihood of re-identification was negligible. On this point, the Commission’s guidelines [40] on the free flow of non-personal data suggest that it is often difficult to assess the effectiveness of an anonymization procedure. Indeed, besides many academic papers [3, 56, 31, 74, 1, 30], even a study commissioned by the European Parliament’s ITRE Committee has shown that it is possible to re-identify supposedly anonymized data [48].
The risk of re-identifying anonymous data stems from the technical limitations of anonymization and the lack of clear and realistic guidelines on the subject. This suggests that re-identification is not only a consequence of poor anonymization implementation.
Moreover, national DPAs disagree on how to implement anonymization. The French DPA, CNIL, adopts the WP29’s approach to anonymization; other national DPAs are more flexible. For example, the UK’s ICO (Information Commissioner’s Office), despite UK’s own data protection regulation being largely similar to that of the EU, states that [60]:
@@
“Organisations don’t have to be able to prove that it is impossible for any data subject to be identified in order for an anonymization technique to be considered successful. Rather, if it can be shown that it is unlikely that a data subject will be identified given the circumstances of the individual case and the state of technology, the data can be considered anonymous.”
The CNIL’s guidelines on anonymization are of particular interest, as they establish a strict standard to determine whether data can be considered anonymous. They assert that data is anonymous only when it is impossible to re-identify the data subject. However, they recognise that when the risks of singling out, linkability, and inference are not met, data can be deemed anonymous if a subsequent analysis indicates a negligible risk of re-identification [23]. Unfortunately, the definition of what constitutes a negligible risk remains ambiguous, since anonymization is a context-dependent process, depending on the nature of the data and its intended use.
Take-away The lack of a clear, harmonized definition of anonymization across EU legal texts and among national authorities aggravates the complexities of anonymization and favors the confusion between pseudonymization and anonymization. Although anonymization and pseudonymization rely on similar techniques, they have different aims: anonymization aims to conceal the data subject’s identity, whereas pseudonymization aims to protect privacy by making re-identification more complex but not impossible. Technical documents should help practitioners assess when to opt for anonymization rather than pseudonymization. However, most of the available technical guidance does not address this issue.
@@ 4.2 Technical Documents
  • DPC (Ireland) stands out with well-structured and clear guidance [33, 34] on legal questions; however, it does not offer practical advice.
@@
Some statistical agencies provide additional resources. The National Institute of Statistics and Economic Studies in France (INSEE) offers slides and working papers [63, 65], but most lack practical detail. A notable exception is Bergeat’s work [70] comparing and explaining experiments done by two anonymization software tools: $\mu$-Argus and SDCMicro. It also gives plenty of citations, however, only to sources on statistics (no computer science references). It is aimed at statisticians, and it could serve as a continuation to other introductory materials, but the reader who already has at least a Bachelor’s degree in mathematics or statistics. Other guides, such as [64], focus on confidentiality rules within the French statistical service rather than techniques. Statistical documents only mention statistical tools and use a different language from that of computer scientists. This could, unfortunately, result in ignoring some state-of-the-art methods, such as differential privacy. A good example is the working paper [32] that details the anonymization process applied to a large French administrative database where the authors experimented with different methods including k-anonymity, all-m anonymity, and l-diversity. They mention that they have tried to apply DP; however, they abandoned the experiment due to a lack of expertise.
@@ 5 Inadequate Practices
EU case law lacks clarity with regard to anonymization practices. The most relevant cases focus on clarifying the concept of personal data. However, the interpretations provided help to assess what an anonymized dataset is not. Moreover, the Court’s application of Recital 26 to real cases provides valuable insight into the question of whether data remain anonymous despite a residual risk of re-identification. The following subsections focus on relevant case law concerning the notion of personal data, the implications of which are important for anonymization, and inadequate anonymization practices, demonstrating a lack of awareness on the subject both at the institutional and organizational levels. The final subsection also addresses the inadequate practices that arise from the confusion between anonymization and pseudonymization. At least one example is given for each subsection; they refer to the most representative and recent cases, but these are not exhaustive of all the existing case law on the subject.
5.1 Case-law on personal data
@@
The General Court’s SRB vs EDPS decision is a good example of a decision relating to personal data whose implications are also relevant for anonymization [54]. SRB (Single Resolution Board) carried out an insolvency procedure against Banco Popular. Within this procedure, some data were processed to assess the eligibility of the participants for compensation. Each participant was identifiable by means of an alphanumeric code generated randomly. The staff processing these data only had access to these codes and not to the key identifying them. The EDPS (European Data Protection Supervisor) considered these data pseudonymized [47], but its decision was challenged before the General Court. Using a risk-based approach, the General Court decided that the data were anonymous. Indeed, according to Breyer’s Court of Justice case law[50], the additional information (the key) needed to re-identify the data subjects remained inaccessible to the processing staff. The fact that the staff could not legally access the complementary data that would allow re-identification proved enough to consider that no reasonable means existed to re-identify the data, which thus remained non-personal [54].
@@
This example shows that the EU Court of justice’s (EUCJ) case law on personal data builds upon its precedents rather than undergoing a radical evolution. On this point, Breyer’s decision [50] was about the dynamic nature of IP addresses. These IP addresses are subject to change with each connection. The plaintiff initiated legal proceedings against the Republic of Germany for its practices concerning the storage and registration of these data. The Court of Justice had to determine whether dynamic IP addresses should be considered personal data for the service provider. The Court decided that the retention of all information by a single individual was not a prerequisite for data being considered as personal. This meant that a third party could retain such re-identifying information, and that this circumstance did not affect the qualification of the data. However, the Court acknowledged that an assessment was necessary to determine the reasonableness of combining this information, taking into account the effort, time, and cost associated with the operation, as well as the accessibility of this additional information (enabling user identification) to the service provider. Given the legal restrictions on such access in Germany, the Court determined that in the absence of legal means to obtain this information, the data in question were not deemed personal. The doctrine posits that two fundamental elements have been applied since the Breyer decision to ascertain the personal nature of data. These are: (1) the distinguishability of the data, defined as the capacity of the data points to identify an individual, and (2) the availability of additional data to “situationally relevant entities” who are capable of associating these data with a physical person[88].
@@
The Scania decision perfectly [53] illustrates this methodology. In this case, the Court was asked to determine the legal status of a vehicle identification number (VIN), a unique alphanumeric code assigned by manufacturers to identify the proprietor of a vehicle. In its decision, the Court stated that the VIN can be personal data for independent operators and vehicle manufacturers if the former have the additional data that enable re-identification, and for the latter if they make the VIN available. The availability of data is considered in conjunction with the capacity of isolating owners of vehicles or all other people who have a title on them. Some authors have suggested an evolution in the interpretation of personal data, attributing this change to the Court’s categorization of independent operators and manufacturers as “situationally relevant entities”, capable of associating VIN with additional identifying information.
@@
Take-away The consistency of the Court’s jurisprudence on the concept of personal data is paramount to contrast the phenomenon of anonymity-washing. This is particularly crucial given the occurrence of poor anonymization or privacy-protecting practices, which the EUCJ is entitled to sanction.
5.2 Case-law on inadequate practices